# Shrinking the Generation-Verification Gap with Weak Verifiers

### v1

@@ B.3 Adaptation Method
    Binarization: We use a small amount of labeled samples $\mathcal{D}^{\text{dev}}$ (which we already are using to compute $\Pr(y=1)$) to determine a threshold for converting continuous verifier outputs to binary outputs. Fig. 13 illustrates the precision matrix of the verifier scores after binarization. It shows a damping of large off-diagonal dependencies and improved condition numbers. Table 4 shows that with only 5 to 10 labeled queries from benchmark development sets (which $\leq$ 1% of the evaluation set), we can estimate binarization thresholds that bolster performance by averages of 8.4% for the 70B models when compared to binary splitting along the median score of the dataset.
@@
    Fig. 14 illustrates the precision matrix of the verifier scores after both binarization and dropping low-signal or redundant verifiers. Compared to Fig. 13, it shows further attenuation of off-diagonal structure. Illustrating that dropping contributes substantially to decorrelating the verifier set, which can improve identifiability and numerical stability for downstream weak supervision. As shown in Table 5, verifier pruning leads to 12.5% performance improvement for the 70B model setting.
@@ B.4 Exploration: Clustering by Difficulty to Improve Weaver
Weak verifiers often behave inconsistently across the difficulty spectrum of input queries. For instance, most verifiers may have very high accuracy on easy queries and low accuracy on more difficult queries. Fig. 11 and Fig. 10 illustrate the average and selection accuracy of each verifier across multiple problems and datasets. This confirms that there is significant variation in verifier performance at the per-query level. To capture this heterogeneity, we explore clustering queries by difficulty and fitting one Weaver’s weak supervision model per cluster, independently.
@@ C.2.2 Alternative Verification Strategies
Logistic Regression: We train a logistic regression classifier where the input features are the verifier scores $[s_{ij1},...,s_{ijm}]$ and the output is the correctness of each response: $P(y_{ij}=1|\mathbf{s}{ij})=\sigma(\mathbf{w}^{T}\mathbf{s}{ij}+b)$, where $\sigma$ is the sigmoid function. The weights $\mathbf{w}$ and bias $b$ are learned using labeled training data. This supervised approach can capture more complex relationships between verifier outputs than naive averaging but requires substantial labeled data for effective training.
@@ C.3 Scaling Trends of Weaver
Scaling laws describe how performance metrics such as accuracy, sample efficiency or compute cost change as we scale controllable resources, i.e. the number of trials $K$, model capacity. [37] showed that, for fixed-parameter Transformer language models, the cross-entropy loss decreases as a power-law in both model size and data. This framework has since been extended to explore optimal tradeoffs between model and data scaling [33], as well as inference-time scaling with multiple samples [13, 6].
@@
$X_{i}=\begin{cases}1,&\text{if the i-th query is solved at least once (with probability $q_{i}$)},\\ 0,&\text{otherwise},\end{cases}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Our goal is to compute the probability that the maximum of $c$ i.i.d draws from $f_{1}$ exceeds the maximum of $K-c$ draws from $f_{0}$.
@@
[…]
@@
[…]
@@
To obtain practical scaling trends, we fit parametric models in Eq. 14 to the empirical averages computed from $5$ independent runs for each value of $K$, across each dataset and verification strategy. Specifically, we use the L-BFGS-B algorithm to optimize a smooth approximation following [32]. To ensure numerical stability and robustness to outliers or heavy-tailed noise in the observed selection accuracies, we minimize the Huber loss between the predicted values and the empirical means. The Huber loss behaves quadratically for small residuals and linearly for large ones, making it less sensitive to outliers than mean squared error (MSE) while maintaining smooth differentiability for gradient-based optimization. It is defined as,
@@
[…]
@@
Fig. 16 and Fig. 17 along with Table 18 illustrate how the different verification strategies scale with the number of generations and the fit to the parametric form in Eq. 14. Each method exhibits characteristic scaling behavior that aligns with Eq. 14. Weaver demonstrates improved performance over naive ensembles and majority voting. The fitted parameters in Table 18 quantitatively capture these trends across datasets, providing evidence that the parametric from in Eq. 15 closely model empirical outcomes. Fig. 18 and Fig. 19 along with Table 19 illustrate the predictive performance of the parametric form in Eq. 15, showing that models fit on subsets of $K$ can extrapolate to unseen values of $K$.
@@
Table 18: Fitted parameters for Scaling Trends in Fig. 16 and Fig. 17.
@@
Table 19: Fitted parameters for Scaling Trends with 90% of data in Fig. 18 and Fig. 19.
@@ C.4 Scaling Candidate Generations
Methodology Generations ($K$) Datasets Average
@@
Baselines First Sample 1 49.8% 28.3% 64.1% 46.6% 47.2%
@@ C.5 Scaling Verifier Count
When breaking down weak verifiers into RMs or LM Judges, individually, we find that additional LMs leads average gains of 5.4% and 6.1%, respectively (Table 21). In contrast, sampling additional scores from a single RM or LM judge yields only 0.8% and 1.1% gains on average. These results suggest that leveraging the complementary strengths of multiple verifiers can be more effective than eliciting multiple judgments from a single verifier. Sec. C.7 provides additional details on the verifier prompting. Finally, Figure 22 illustrates the tradeoff of scaling the number of verifiers versus increasing the number of scores from a single verifier, showing that scaling verifiers is helpful when the coverage increases as we increase sample count.
@@
In Fig. 22 illustrates how the number of verifiers and repeated generations interact to influence success rate. We observe that increasing the number of generations tends to be more effective than increasing the number of verifiers alone—but only when paired with the right verification strategy. For example, naive ensembling of verifiers plateaus in performance even as more generations are added, whereas Weaver continues to improve with both axes. This highlights that generation diversity is a stronger driver of performance than verifier count alone, and that weak supervision methods like Weaver are essential to fully leverage this diversity. We illustrate the verification generation tradeoff for additional datasets in Appendix C.3.
@@ C.7.1 LM Judge Prompt Optimization
[…]

### v2

@@ B.3 Adaptation Method
    Binarization: We use a small amount of labeled samples $\mathcal{D}^{\text{dev}}$ (which we already are using to compute $\Pr(y=1)$) to determine a threshold for converting continuous verifier outputs to binary outputs. Fig.˜13 illustrates the precision matrix of the verifier scores after binarization. It shows a damping of large off-diagonal dependencies and improved condition numbers. Table 4 shows that with only 5 to 10 labeled queries from benchmark development sets (which $\leq$ 1% of the evaluation set), we can estimate binarization thresholds that bolster performance by averages of 8.4% for the 70B models when compared to binary splitting along the median score of the dataset.
@@
    Fig.˜14 illustrates the precision matrix of the verifier scores after both binarization and dropping low-signal or redundant verifiers. Compared to Fig.˜13, it shows further attenuation of off-diagonal structure. Illustrating that dropping contributes substantially to decorrelating the verifier set, which can improve identifiability and numerical stability for downstream weak supervision. As shown in Table 5, verifier pruning leads to 12.5% performance improvement for the 70B model setting.
@@ B.4 Exploration: Clustering by Difficulty to Improve Weaver
Weak verifiers often behave inconsistently across the difficulty spectrum of input queries. For instance, most verifiers may have very high accuracy on easy queries and low accuracy on more difficult queries. Fig.˜11 and Fig.˜10 illustrate the average and selection accuracy of each verifier across multiple problems and datasets. This confirms that there is significant variation in verifier performance at the per-query level. To capture this heterogeneity, we explore clustering queries by difficulty and fitting one Weaver’s weak supervision model per cluster, independently.
@@ C.2.2 Alternative Verification Strategies
Logistic Regression: We train a logistic regression classifier where the input features are the verifier scores $s_{ij1},...,s_{ijm}]$ and the output is the correctness of each response: $P(y_{ij}=1|\mathbf{s}{ij})=\sigma(\mathbf{w}^{T}\mathbf{s}{ij}+b)$, where $\sigma$ is the sigmoid function. The weights $\mathbf{w}$ and bias $b$ are learned using labeled training data. This supervised approach can capture more complex relationships between verifier outputs than naive averaging but requires substantial labeled data for effective training.
@@ C.3 Scaling Trends of Weaver
Scaling laws describe how performance metrics such as accuracy, sample efficiency or compute cost change as we scale controllable resources, i.e. the number of trials $K$, model capacity. [37] showed that, for fixed‐parameter Transformer language models, the cross-entropy loss decreases as a power-law in both model size and data. This framework has since been extended to explore optimal tradeoffs between model and data scaling [33], as well as inference-time scaling with multiple samples [13, 6].
@@
[…]
@@
The expected coverage $(\text{Pass@K})$ is the expected fraction of problems solved after trying $K$ times per problem:
$\text{Pass@K}:=\mathbb{E}[Y]/N=\frac{1}{N}\sum_{i=1}^{N}(1-(1-p_{i})^{K})$
@@
$\log[-\log\text{Pass@K}]$ $=\log\Gamma(\beta+K)+\log\Gamma(\alpha+\beta)-\log\Gamma(\beta)-\log\Gamma(\alpha+\beta+K)$
@@
$\text{Selection@1}(K):=\mathbb{P}[\text{top-scoring response is correct}]$ (13)
@@
$\text{Selection@1}=\mathbb{P}\left[\max s^{(1)}>\max s^{(0)}\right]$
Our goal is to compute the probability that the maximum of $c$ i.i.d draws from $f_{1}$ exceeds the maximum of KcK-c draws from $f_{0}$.
@@
$\text{Selection@1}(K)=\mathbb{E}_{p\sim\text{Beta}(\alpha,\beta)}\left[\text{Selection@1}(K\mid p)\right]$
@@
$\text{Selection@1}(K)\approx\exp(-\zeta K^{-\alpha})\cdot\left(1-(1-\pi)^{K^{\gamma}}\right)$ (14)
@@
To obtain practical scaling trends, we fit parametric models in Eq.˜14 to the empirical averages computed from $5$ independent runs for each value of $K$, across each dataset and verification strategy. Specifically, we use the L-BFGS-B algorithm to optimize a smooth approximation following [32]. To ensure numerical stability and robustness to outliers or heavy-tailed noise in the observed selection accuracies, we minimize the Huber loss between the predicted values and the empirical means. The Huber loss behaves quadratically for small residuals and linearly for large ones, making it less sensitive to outliers than mean squared error (MSE) while maintaining smooth differentiability for gradient-based optimization. It is defined as,
@@
$\text{Selection@1}(K)\approx\text{floor}+(\text{ceil}-\text{floor})\cdot\exp(-\zeta K^{-\alpha})\cdot\left(1-(1-\pi)^{K^{\gamma}}\right)$ (15)
@@
Fig.˜16 and Fig.˜17 along with Table˜18 illustrate how the different verification strategies scale with the number of generations and the fit to the parametric form in Eq.˜14. Each method exhibits characteristic scaling behavior that aligns with Eq.˜14. Weaver demonstrates improved performance over naive ensembles and majority voting. The fitted parameters in Table˜18 quantitatively capture these trends across datasets, providing evidence that the parametric from in Eq.˜15 closely model empirical outcomes. Fig.˜18 and Fig.˜19 along with Table˜19 illustrate the predictive performance of the parametric form in Eq.˜15, showing that models fit on subsets of $K$ can extrapolate to unseen values of $K$.
@@
Table 18: Fitted parameters for Scaling Trends in Fig.˜16 and Fig.˜17.
@@
Table 19: Fitted parameters for Scaling Trends with 90% of data in Fig.˜18 and Fig.˜19.
@@ C.4 Scaling Candidate Generations
Methodology
Generations ($K$)
Datasets
Average
@@
Baselines
First Sample 1 49.8% 28.3% 64.1% 46.6% 47.2%
@@ C.5 Scaling Verifier Count
When breaking down weak verifiers into RMs or LM Judges, individually, we find that additional LMs leads average gains of 5.4% and 6.1%, respectively (Table 21). In contrast, sampling additional scores from a single RM or LM judge yields only 0.8% and 1.1% gains on average. These results suggest that leveraging the complementary strengths of multiple verifiers can be more effective than eliciting multiple judgments from a single verifier. Sec.˜C.7 provides additional details on the verifier prompting. Finally, Figure 22 illustrates the tradeoff of scaling the number of verifiers versus increasing the number of scores from a single verifier, showing that scaling verifiers is helpful when the coverage increases as we increase sample count.
@@
In Fig.˜22 illustrates how the number of verifiers and repeated generations interact to influence success rate. We observe that increasing the number of generations tends to be more effective than increasing the number of verifiers alone—but only when paired with the right verification strategy. For example, naive ensembling of verifiers plateaus in performance even as more generations are added, whereas Weaver continues to improve with both axes. This highlights that generation diversity is a stronger driver of performance than verifier count alone, and that weak supervision methods like Weaver are essential to fully leverage this diversity. We illustrate the verification generation tradeoff for additional datasets in Appendix C.3.
@@ C.7.1 LM Judge Prompt Optimization
Our experimental setup uses training examples containing instruction-generation pairs. Since our datasets have multiple generations per instruction (up to 100), we group examples by instruction before splitting to prevent data leakage between train and validation sets. We hold out 50% of the dataset instructions (each paired with 100 candidate generations) for evaluation. For the optimization space scaling experiment, we randomly select $n$ generations such that $n\times\text{len(dataset)}/2=250$, maximizing training set diversity while maintaining a fixed training set size. For the data scaling experiment, we train on different percentages (1%, 2%, 4%, and 16%) of the dataset by calculating the number of instructions as $\lceil\text{num\_problems\_in\_dataset}\times(\text{train\_percentage}/100)\rceil$ and selecting repeated samples for each instruction with $\text{samples}=\min(\max(4,\text{num\_problems}\times 2),20)$ to avoid overfitting. We use a consistent random seed to ensure identical dataset splits between optimization runs.